find the equation of the quadratic function whose graph has the vertex (0,2) and contains the point (1,3)
Answer by LinnW(1048)

Since the vertex is (0,2)
our equation has the form y = ax^2 + 2
using the point (1,3) for (x,y)
y = ax^2 + 2 becomes 3 = a*1^2 +2
3 = a*1 + 2
3 = a + 2
add -2 to each side
1 = a
So our equation is y = 1*x^2 +2
or y = x^2 + 2